OpenStudy (anonymous):

Solve using algebraic methods. y=x squared -4x+6 and y=2

OpenStudy (anonymous):

\[y=x^2-4x+6\]\[2=x^2-4x+6\]\[0=x^2-4x+4\]\[0=(x-2)(x-2)\]\[x=2\]